1) Define impulse and momentunm 2) A 1500kg car is traveling at 20m/s and the driver notices a deer jump in front of the car. The driver breaks and the car slows down to 15 m/s as the car hits the deer .5s later. What is the total force the car exerts onto the deer on impact?

explain in details and calculations please
Answer #1

1)

Impulse

It is defined as the product of force and time


Let 'F' be the force acts on a body for a period of time 't' then the impulse is given by


I = F * t


Momentum


It is defined as the product of mass and velocity

Let a body of mass 'm' moves with a velocity 'v' then the momentum is

p = m * v

2)


Given

mass of car is m = 1500 kg

initial velocity v1 = 20 m/s

After applying brakes final velocity v2 = 15 m/s

The time for which the force is applied is t = 5 s


Let 'F' be the total force


From impulse momentum theorem


Impulse = change in momentum

I = p2 - p1

F * t = (m * v2) - (m * v1)

F * t = m * (v2 - v1)

F * 5 = 1500 * (15 - 20)

F = - 1500 N

The negative sign says the applied force decreases the speed


The total force the car exerts onto the deer on impact is   1500 N